OpenStudy (anonymous):

I have no idea For two quadrilaterals A B C D is similar to Q R S T, line segment A B equals 3 inches, line segment B C equals line segment C D equals 4 inches, and line segment D A equals 6 inches. Match the given scale factor with the given side length between quadrilaterals A B C D and Q R S T. Match Term Definition 2 A) line segment T Q equals 3 inches zero point five B) line segment S T equals 6 inches 3 C) line segment R S equals eight inches one point five D) line segment Q R equals 9 inches

OpenStudy (anonymous):

@cwrw238

OpenStudy (gorv):



OpenStudy (gorv):

DID U MISSED THE LENGTH OF BC??

OpenStudy (anonymous):

whaaaaaaaaaat?

OpenStudy (gorv):

line segment BC equal ...in Q after this any value is there ??? match the Q u have and u typed plzz

OpenStudy (gorv):

for similar quadrilateral

OpenStudy (gorv):

\[\frac{ AB }{ QR }=\frac{ BC }{ RS }=\frac{ CD }{ ST }=\frac{ DA }{ TQ }\]

OpenStudy (gorv):

ratio of corresponding side will be equal

OpenStudy (gorv):

\[\frac{ AB }{ BC }=\frac{ QR }{ RS }\]

OpenStudy (gorv):

\[\frac{ 3 }{ 4 }=\frac{ QR }{ RS }\]

OpenStudy (gorv):

u checked ????????????????????????????????????????????

OpenStudy (gorv):

but QR=9 RS=8